OpenStudy (anonymous):

Kamala spends 20% of her monthly salary for her son's education. When she will get her increment of Rs.170 next month she plans to spend half of that also for her son's education. The total she would spend then, would be Rs.645. What is her present salary?

OpenStudy (anonymous):

@ParthKohli pls help

OpenStudy (anonymous):

@felixsalter can u help

OpenStudy (anonymous):

@ash2326 Please help

OpenStudy (anonymous):

Please some one help

OpenStudy (ash2326):

Let her current salary be Rs. x then she spends 20 % of x on her son' s education= \(\frac{20}{100}\times x=Rs\ 0.2x\) Now she gets an increment of 170 next month, she will spend half of this ( 170/2=85) also on her son's education so total she will spend =0.2x+85 and this is given as 645 \[0.2x+85=645\] From this you'll find x which is her salary Can you do this now?

OpenStudy (anonymous):

thanx got it